OpenStudy (anonymous):

A dollhouse uses the scale 1:10. A chair in the dollhouse is 8 centimeters tall. How tall is the actual chair? 8 m 0.08 m 0.8 m 80 m

jimthompson5910 (jim_thompson5910):

let h = height of the chair So if the scale is 1:10 and the chair in the dollhouse is 8 cm, then 1/10 = 8/h Solve this for h and tell me what you get

OpenStudy (anonymous):

do you solve h by dividing1/10

OpenStudy (anonymous):

do i set up a proportion

jimthompson5910 (jim_thompson5910):

you can, but i find it easier to cross multiply 1/10 = 8/h 1*h = 10*8 h = 80 Since h is in cm, this means 80 cm = 0.8 m

OpenStudy (anonymous):

o

jimthompson5910 (jim_thompson5910):

since (80 cm)*(1 m/100 cm) = 80/100 = 0.8 m

OpenStudy (anonymous):

gotcha